# On-device Collaborative Language Modeling via a Mixture of Generalists and Specialists

### v1

@@ 4.2.1 Result Analysis
The comparison between our method and the baseline methods is summarized in Table 1. We provide an extension of baseline comparison in Appendix C.1.
@@
Depending on the dataset, either CoMiGS-2G or CoMiGS-2S achieves the highest performance. The key distinction compared to Local or FedAvg is the existence of a layer-wise router, which weighs the two generalists or two specialists for each token, as opposed to assigning equal weights. This highlights the effectiveness of our learned routing mechanism.
Token-level vs. client-level collaborative decisions:
Compared to the state-of-the-art baseline, PCL, as proposed by Wagner et al. (2024), our method demonstrates a clear performance improvement. PCL assigns a pairwise collaboration weight between users by evaluating how well user $i$’s model performs on user $j$’s validation set. On the two in-domain tasks, PCL exhibits performance similar to Local, where the resulting collaboration matrices are nearly identity matrices, thereby limiting effective collaboration between users. Our method, in contrast, decides the collaboration pattern based on each input token, and thus can harness the collective power more effectively.
The impact of co-existence of generalists and specialists:
The performances of CoMiGS-2G and CoMiGS-2S are not consistent across the different scenarios, while our CoMiGS-1G1S can always closely track the best-performing model, which is clearly shown in Figure 3. Depending on the task type, generalists and specialists alone may not be sufficient. A balanced combination of personalization and collaboration is required, and our approach achieves this balance effectively.
AG News Multilingual SlimPajama
Pretrained 90.65 156.12 37.19
Centralized 28.19 (0.52) 55.41 (0.12) 19.53 (0.14)
Local 41.50 (0.05) 55.49 (0.64) 27.31 (0.17)
FedAvg 32.24 (0.05) 56.90 (0.16) 23.30 (0.17)
PCL 32.11 (0.04) 54.02 (0.19) 27.12 (0.33)
CoMiGS - 2S 35.81 (0.13) 46.36 (0.16) 22.51 (0.08)
CoMiGS - 2G 31.18 (0.05) 58.31 (0.17) 21.36 (0.01)
CoMiGS - 1G1S 33.53 (0.03) 47.19 (0.10) 21.79 (0.04)
Table 1: Mean test perplexity over the users with homogeneous models, averaged across 3 seeds. Mean (std) with a rank locator for the mean (the lower the better). Green denotes the best performing methods and red denotes our method.
@@ 4.2.2 Routing Analysis
Figure 4 illustrates the evolution of averaged layer-wise routing scores for the generalist and specialist experts. As training progresses, specialists gradually take over the role of generalists across all layers. Notably, for out-of-domain tasks such as AG News, generalists play a more significant role early in training. By the end of training, the score distribution stabilizes, with generalists still contributing 30%-50% to next-token predictions, underscoring their importance in handling out-of-domain tasks. For in-domain tasks, more weight is shifted toward specialists. However, generalists still hold considerable importance, especially in datasets like SlimPajama, where some users are locally exposed to math or code-related texts, and generalists are needed to better process standard texts. Additionally, it is observed that deep layers have more skewed score distributions, while shallow layers weigh generalists and specialists more evenly.
Token-wise analysis:
We further present a token-level routing result visualization on models fine-tuned with SlimPajama dataset in Figure 5: Users 0 and 1 are fine-tuned with very specific math and programming texts, and they tend to utilize the generalist more in the last layer. Function words (”and”, ”a”, ”on” ”the” etc) are more routed to generalists, as expected. This can be seen in the top right panel of Figure 5. It is important to note that only the top choice is highlighted here. The abundance of blue does not imply that generalist experts play no role in predicting the next token. To see this, compared to when only specialists are present (CoMiGS-2S), our CoMiGS-1G1S gives more consistent results. More detailed token-wise routing result visualization including out-of-domain tasks can be seen in Appendix E. When dealing with out-of-domain texts, there is an increasing tendency to seek for generalists, as shown in the off-diagonal entries in Fig. 15.
Figure 4: Expert Scores for the generalist expert and the specialist expert, averaged across all tokens and multiple batches for all datasets. Darker colors represent deeper layers.
Figure 5: Visualization of in-domain token-level routing results for 1G1S trained on SlimPajama. Tokens are colored with the Top1 expert choice at the first layer (top) and last layer (bottom). Orange denotes the generalist and blue denotes the specialist. Users 0-3 are trained with the categories StackExchange, GitHub, ArXiv, and Book respectively. Texts are generated by ChatGPT.
@@ 4.3.1 Baseline Comparison
In this section, our focus is to deal with system heterogeneity, where users can have different numbers of experts. We still keep one generalist expert, but the number of specialists can vary across the users (our method is denoted as CoMiGS-1GXS). It’s important to note that the richness of computational resources doesn’t always correlate with the complexity of local data. For instance, some users may have ample computational resources but local data in small quantities. In such cases, a crucial objective is to prevent overfitting due to redundant model-fitting abilities.
We compare our approach to two state-of-the-art baselines: HetLoRA from Cho et al. (2023) and FlexLoRA from Bai et al. (2024), both of which adapt LoRA ranks based on the resource capacity of each user. HetLoRA aggregates LoRA matrices ${\bm{A}}$ and ${\bm{B}}$ by zero-padding to the maximum rank and then distributes them back using rank truncation. In contrast, FlexLoRA first reconstructs model updates $\Delta{\bm{W}}$ and redistributes the aggregated updates using SVD. We compare our method to these baselines by matching the number of fine-tunable parameters, measured as both active and full parameters. For example, to match the full parameter count of 1GXS with $(4,2,2,2)$ LoRA experts (rank 8), LoRA modules of ranks $(32,16,16,16)$ would be required. With Top2 routing, to match the active parameter count, each user would need LoRA modules of rank 16.
Our results, shown in Table 2, are based on randomly assigning varying capabilities to users, independent of their local data size, as indicated by the number of LoRA modules per user. We find that our method outperforms the baseline methods most of the time, regardless of whether we match the full parameter count or the active parameter count. This advantage stems from the fact that both HetLoRA and FlexLoRA average model parameters across users without allocating parameters for local adaptations, focusing on building a strong generalist model. In contrast, our approach adaptively integrates both generalist and specialist knowledge, excelling in scenarios where specialized knowledge is crucial.
@@
Active Full Active Full
AG News
(4,2,2,2) 33.66 (0.07) 31.58 (0.14) 31.95 (0.13) 36.45 (0.06) 36.49 (0.17)
(2,4,4,4) 34.22 (0.09) 31.58 (0.14) 32.52 (0.19) 36.45 (0.06) 36.40 (0.08)
Multilingual
@@
Table 2: Mean test perplexity (std) over users with heterogeneous models, averaged across 3 seeds, with red being the top1 method. For example, $(4,2,2,2)$ means in our 1GXS setup users have 4, 2, 2, and 2 experts, respectively, and in the two baselines, all users have rank 16 to match active parameter count, or ranks 32, 16, 16, and 16 to match full parameter count.
@@ 4.3.2 Analysis related to local data quantities
In this section, we further separate resource abundance from data quantity. Our approach is more robust to overfitting due to the regularizing effect of the generalist, while at the same time better fitting local data through the incorporation of specialist knowledge.
@@
From Figure 3 we observe that the local data complexity for User 0 and User 1 is quite low: doing local model training with two times LoRA modules is prone to over-fitting. This is confirmed by low data quantity in both users in Figure 2. Instead, User 2 and User 3 benefit from longer local training. In the experiment shown in Figure 6 we validate that providing User 2 and User 3 more resources, by increasing the number of local experts to 4 and 8 (while still keeping top2 routing) indeed increases performance for both users.
Figure 6: Test Perplexity during training for the multilingual setup. User 2 and User 3 have more data locally and thus benefit from having more experts. The numbers in the legend indicate the number of experts $n_{i}$ within each user. Top-2 routing is performed.
@@
In practice, users may not know their local data complexity, leading to a potential mismatch in resource allocation relative to data quantity. To simulate such scenarios, we allocate more experts to users with less data, specifically User 0 and User 1. Figure 7 demonstrates that our method is resilient to overfitting, even when fine-tuning twice or four times as many expert parameters.
Figure 7: In the multilingual setup, Users 0 and 1 despite having high resources locally, do not overfit on their small-sized local data. Numbers in the legend denote the number of experts $n_{i}$ within each user. Top2 routing is performed.
@@
What if users can only support at most one expert? When those users participate in collaboration, in our setup they have to go with the generalist expert. We show that if we allocate more specialists to users with a higher local data quantity, the generalist knowledge can further be refined, as illustrated by the performance of User 0 and User 1 in Figure 8.
Figure 8: In the multilingual setup, Users 0 and 1, despite having only one expert locally, still benefit from other users having more experts, thereby enhancing the generalist’s performance. The numbers in the legend indicate the number of experts, $n_{i}$, within each user. Top-2 routing is applied when $n_{i}\geq 2$.
@@ 5 Conclusions and Future Directions
Building on the Mixture of Experts (MoE) architecture, we propose a novel framework for on-device personalized collaborative fine-tuning of LLMs, which combines generalist and specialist experts. Our carefully designed routing mechanism ensures a balance between personalization and collaboration for each input token. This data-dependent approach consistently delivers strong performance across various datasets and task distributions.
Moreover, our framework adapts to varying local data sizes and computational resource constraints. High-resourced users can leverage larger datasets for enhanced performance while maintaining resilience against overfitting in scenarios with limited data.
@@ Appendix A Limitations and Societal Impact
We offer a collaboration framework for edge devices, aiming to enable smaller devices to leverage large language models (LLMs) despite limited resources and data availability. Our approach enhances fairness and mitigates privacy concerns by ensuring data remains on end devices. The privacy aspects can further be enhanced by differential privacy, which we do not pursue here.
@@ Appendix B Extra Experimental Details
The number of tokens within each user is shown in Table 3.
user 0 user 1 user 2 user 3
AG News Training 761’924 756’719 814’131 771’460
Validation 48’809 48’730 50’398 48’249
Test 48’167 47’721 48’344 49’377
@@ Appendix C More Tables and Figures
C.1 Extended Baseline Comparison
@@
AG News Multilingual SlimPajama
@@
Pretrained 90.65 156.12 37.19
Centralized 28.19 (0.52) 55.41 (0.12) 19.53 (0.14)
Local 41.50 (0.05) 55.49 (0.64) 27.31 (0.17)
FedAvg 32.24 (0.05) 56.90 (0.16) 23.30 (0.17)
FFA-LoRA 31.60 (0.14) 57.83 (0.13) 23.42 (0.069)
PCL 32.11 (0.04) 54.02 (0.19) 27.12 (0.33)
@@
CoMiGS - 2S 35.81 (0.13) 46.36 (0.16) 22.51 (0.08)
CoMiGS - 2G 31.18 (0.05) 58.31 (0.17) 21.36 (0.01)
CoMiGS - 1G1S 33.53 (0.03) 47.19 (0.10) 21.79 (0.04)
@@ C.2 HetLoRA
C.3 Is the standard load balancing loss sufficient?
The standard load balancing loss encourages equal assignment of tokens to each expert. To encourage enough tokens to be routed to the generalist expert such that more general knowledge can be developed, we add a modified load-balancing loss to the outer optimization objective in eq equation 4 by introducing an importance weighting. We separate the 0-th expert to be the generalist expert and conduct Top-2 routing. The modified load balancing loss is as follows:
@@ C.3 Is the standard load balancing loss sufficient?
$j$ is the expert index and $p(x)=[p_{j}(x)]_{j=1}^{n_{i}}$ is the logit output from the routing network for a specific token $x$. The idea is that one of the top 2 tokens should always be routed to the generalist expert, i.e. the 0-th expert. Thus, $\frac{p_{0}}{1/2}$ should be equal to $\frac{p_{i}}{1/2(n_{i}-1)}$ for $i\neq 0$. As the original load balancing loss encourages uniform distribution, this modification encourages the generalist expert to have a routing probability of 0.5 on expectation. Note that when $n_{i}=2$, this ${\mathcal{L}}_{i}^{\text{LB}}$ is the same as the original load balancing loss as proposed in Fedus et al. (2022).
@@ Appendix D Additional Experiments
We replicate the experiments in Section 4.3 with the SlimPajama dataset, where we assign four times as many tokens to Users 2 and 3 as to Users 0 and 1.
More Specialists Help with Higher Data Quantity. From Figure 10, it is clear that users 2 and 3 benefit from more experts locally.
Figure 10: Test Perplexity during training for the SlimPajama setup. User 2 and User 3 have more local data and thus benefit from having more experts. The numbers in the legend indicate the number of experts $n_{i}$ within each user. Top-2 routing is performed.
@@
Figure 11: In this SlimPajama setup, Users 0 and 1 despite having low resources locally, overfit slightly on their small-sized local data. Numbers in the legend denote the number of experts $n_{i}$ within each user. Top2 routing is performed.
Specialists Can Benefit Generalists. Low-resourced users that can only support a single expert setup still benefit from collaboration, as the generalist knowledge is refined through a more detailed distinction between specialist and generalist roles via other high-resourced users.
Figure 12: In this SlimPajama setup, Users 0 and 1, despite having only one expert locally, still benefit from other users having more experts, thereby enhancing the generalist’s performance. The numbers in the legend indicate the number of experts, $n_{i}$, within each user. Top-2 routing is applied when $n_{i}$geq 2

### v2

@@ 4.2.1 Result Analysis
The comparison between our method and the baseline methods is summarized in Table 1.
@@
Depending on the dataset, either CoMiGS-2G or CoMiGS-2S achieves the highest performance. The key distinction compared to Local or FedAvg is the existence of a layer-wise router, which weighs the two generalists or two specialists for each token according to the validation performances, as opposed to assigning equal weights. This emphasizes that even with the same expert knowledge, the way it’s combined is crucial. Moreover, pFedMoE, despite having a learned router as well, underperforms our method, even in the in-distribution scenario. The reason is that the routing parameters are updated simultaneously with the expert parameters using the training set, and thus cannot effectively adapt to the target distribution.
Token-level collaborative decisions outperform Client-level:
Compared to the state-of-the-art baseline PCL, as proposed by Wagner et al. (2024), our method demonstrates a clear performance improvement. PCL assigns a pairwise collaboration weight between users by evaluating how well user $i$’s model performs on user $j$’s validation set. On the two in-distribution tasks, PCL exhibits performance similar to Local, where the learned collaboration matrices are nearly identity matrices, thereby limiting effective collaboration between users. Our method, in contrast, decides the collaboration pattern based on each input token, and thus can harness the collective power more effectively.
The necessity of the co-existence of generalists and specialists:
The performances of CoMiGS-2G and CoMiGS-2S are not consistent across the different scenarios, while our CoMiGS-1G1S can always closely track the best-performing model, which is clearly shown in Table 1 and visualized in Figure 9. Depending on the task type, generalists and specialists alone may not be sufficient. A balanced combination of personalization and collaboration is required, and our approach achieves this balance effectively.
@@ 4.2.2 Routing Analysis
Token-wise analysis:
We further present a token-level routing result visualization on models fine-tuned with SlimPajama dataset in Figure 2: The first two users are fine-tuned with very specific math and programming texts, and they tend to utilize the generalist more in the last layer. Function words (”and”, ”a”, ”on” ”the” etc) are more routed to generalists, as expected. This can be seen in the top right panel of Figure 2. It is important to note that only the top choice is highlighted here. The abundance of blue does not imply that generalist experts play no role in predicting the next token. To see this, compared to when only specialists are present (CoMiGS-2S), our CoMiGS-1G1S gives more consistent results. More detailed token-wise routing result visualization including out-of-distribution tasks can be seen in Appendix E. When dealing with out-of-distribution texts, there is an increasing tendency to seek for generalists, as shown in the off-diagonal entries in Figure 14-19.
@@
Figure 3 depicts the evolution of averaged layer-wise router outputs for the generalist and specialist experts on the out-of-distribution task, comparing CoMiGS-1G1S and pFedMoE. As training progresses, CoMiGS-1G1S undergoes a phase transition: the layer-wise routers initially favor generalists but gradually shift towards specialists. This shift is not observed in pFedMoE, highlighting the critical role of our routing mechanism in handling out-of-distribution tasks. Additionally, we notice different layers converge to a different expert score distribution. When applying our CoMiGS-1G1S, for each user, there are always certain layers where the routers consistently prefer generalists, which aligns with the fact that our target distribution is a union of all local training distributions. This phenomenon no longer occurs with in-distribution tasks, as shown in Figure 10.
Figure 2: Visualization of in-distribution token-level routing results for CoMiGS-1G1S trained on SlimPajama. Tokens are colored with the Top1 expert choice at the first layer (top) and last layer (bottom). Orange denotes the generalist and blue denotes the specialist. Texts are generated by ChatGPT. Further colored text plots are provided in Appendix E.
Figure 3: Expert Scores for the generalist expert and the specialist expert, averaged across all tokens and multiple batches for the out-of-distribution task (AG News), with x-axis being the number of iterations. Upper row: our CoMiGS-1G1S, bottom row: pFedMoE. Darker colors represent deeper layers. Expert score plots for in-distribution tasks can be seen in Figure 10 .
@@ 4.3.1 Baseline Comparison
In this section, our focus is to deal with system heterogeneity, where users can have different numbers of experts. We still keep one generalist expert, but the number of specialists can vary across the users (our method is denoted as One-Generalist-X-Specialists, named CoMiGS-1GXS). It’s important to note that the richness of computational resources doesn’t always correlate with the complexity of local data. For instance, some users may have ample computational resources but local data in small quantities. In such cases, a crucial objective is to prevent overfitting due to redundant model-fitting abilities.
We compare our approach to two state-of-the-art baselines: HetLoRA from Cho et al. (2023) and FlexLoRA from Bai et al. (2024), both of which adapt LoRA ranks based on the resource capacity of each user. HetLoRA aggregates LoRA matrices ${\bm{A}}$ and ${\bm{B}}$ by zero-padding to the maximum rank and then distributes them back using rank truncation. In contrast, FlexLoRA first reconstructs model updates $\Delta{\bm{W}}$ and redistributes the aggregated updates using SVD. We compare our method to these baselines by matching the number of fine-tunable parameters, measured as both active and full parameters. For example, to match the full parameter count of CoMiGS-1GXS with $(4,2,2,2)$ LoRA experts (rank 8), LoRA modules of ranks $(32,16,16,16)$ would be required. With Top2 routing, to match the active parameter count, each user would need LoRA modules of rank 16.
Our results, presented in Table 2, are based on allocating different computational resources to users, with resource availability decoupled from local task complexity. We find that our method outperforms the baseline methods most of the time, regardless of whether we match the full parameter count or the active parameter count. This advantage stems from the fact that both HetLoRA and FlexLoRA average model parameters across users without allocating parameters for local adaptations, focusing on building a strong generalist model. In contrast, our approach adaptively integrates both generalist and specialist knowledge, excelling in scenarios where specialized knowledge is crucial.
@@
CoMiGS-1GXS Active Full Active Full
In Distribution Multilingual
@@
Out of Distribution AG News
(4,2,2,2) 33.66 (0.07) 31.58 (0.14) 31.95 (0.13) 36.45 (0.06) 36.49 (0.17)
(2,4,4,4) 34.22 (0.09) 31.58 (0.14) 32.52 (0.19) 36.45 (0.06) 36.40 (0.08)
Table 2: Mean test perplexity (std) over users with heterogeneous models, averaged across 3 seeds, with red being the top1 method. For example, $(4,2,2,2)$ means in our CoMiGS-1GXS setup users have 4, 2, 2, and 2 experts, respectively, and in the two baselines, all users have rank 16 to match active parameter count, or ranks 32, 16, 16, and 16 to match full parameter count.
@@ 4.3.2 Analysis related to local data quantities
In this section, we further separate resource abundance from data quantity. It is observed that our approach is more robust to overfitting due to the regularizing effect of the generalist, while at the same time better fitting local data through the incorporation of specialist knowledge.
We conduct experiments using Multilingual Wikipedia dataset, where we allocate low data quantities to German and Dutch users, and high data quantities to French and Italian users, as shown in Figure 8. In practice, users may not know their local data complexity, leading to a potential mismatch in resource allocation relative to data quantity. To simulate such scenarios, we allocate model capabilities—measured by the number of LoRA modules per user—either positively or negatively correlated with their local data size. It is important to note that one generalist is always assigned, and resource abundance is only reflected in the number of specialists.
@@
French and Italian users consistently benefit from having more specialists locally, as their test perplexities decrease when the number of specialists increases from 1 to 3 to 7. This suggests that when sufficient local training data is available, adding more specialists leads to improved performance.
Figure 4: Test Perplexity vs. the number of iterations. Low and high denote the relative data quantity among users. The numbers in the legend indicate the number of experts $n_{i}$ within each user. Top-2 routing is performed.
@@
For users with low data quantities, local model training with just two LoRA modules already results in overfitting (a trend observed in Figure 9). Our goal here is to prevent overfitting. Figure 5 demonstrates that our method succeeds to surpress overfitting, even when fine-tuning twice or four times as many expert parameters. We attribute this to the existence of the generalists.
Figure 5: Test Perplexity vs. the number of iterations. Low and high denote the relative data quantity among users. The numbers in the legend indicate the number of experts $n_{i}$ within each user. Top-2 routing is performed. German and Dutch Users despite having high resources locally, do not overfit on their small-sized local data.
@@
What happens if users can only support a maximum of one expert? In our setup, such users must rely on the generalist expert when participating in collaboration. Interestingly, even when their collaborators are allocated more specialists, low-resourced users with only one generalist still benefit from the refined role diversification between generalists and specialists. As a result, the generalists become more powerful, as demonstrated in Figure 6.
Figure 6: Test Perplexity vs. the number of iterations. German and Dutch Users, despite having only one expert locally, still benefit from their collaborators having more experts, thereby enhancing the generalist’s performance. The numbers in the legend indicate the number of experts, $n_{i}$, within each user. Top-2 routing is applied when $n_{i}\geq 2$.
@@ 5 Conclusions and Future Directions
We propose a novel framework for on-device personalized collaborative fine-tuning of LLMs, grounded in an innovative bi-level formulation of the Mixture-of-Experts learning objective. Our fine-grained integration of generalist and specialist expert knowledge achieves superior performance in balancing personalization and collaboration within Federated LLMs.
Furthermore, our framework is the first to address both system and data heterogeneity in collaborative LLM training. It also decouples local data quantity from resource availability, allowing high-resourced users to leverage larger datasets for improved performance while remaining resilient against overfitting in low-data scenarios.
@@ Appendix A Limitations and Societal Impact
We offer a collaboration framework for edge devices, aiming to enable smaller devices to leverage large language models (LLMs) despite limited resources and data availability. Our approach enhances fairness and mitigates privacy concerns by ensuring data remains on end devices. The privacy aspects can further be enhanced by differential private aggregation of generalist weights, which we do not pursue here.
@@ Appendix B Extra Experimental Details
The dataset distribution and number of tokens within each user are shown in Figure 8 and Table 3 respectively.
User 1 User 2 User 3 User 4
@@ Appendix C More Tables and Figures
C.1 Learning curves of different methods
Figure 10: Expert Scores for the generalist expert and the specialist expert from our CoMiGS-1G1S method, averaged across all tokens and multiple batches for the in-distribution task, with x-axis being the number of iterations. Darker colors represent deeper layers.
C.2 Extended Baseline Comparison
@@
In Distribution Out of Distribution
Multilingual SlimPajama AG News
@@
Pretrained 156.12 37.19 90.65
Centralized 55.41 (0.12) 19.53 (0.14) 28.19 (0.52)
Local 54.38 (0.32) 26.95 (0.14) 41.46 (0.06)
FedAvg 58.80 (0.34) 23.27 (0.05) 31.84 (0.02)
FFA-LoRA 57.83 (0.13) 23.42 (0.069) 31.60 (0.14)
PCL 54.53 (0.19) 26.99 (0.19) 32.25 (0.12)
@@
CoMiGS - 2S 46.36 (0.16) 22.51 (0.08) 35.81 (0.13)
CoMiGS - 2G 58.31 (0.17) 21.36 (0.01) 31.18 (0.05)
CoMiGS - 1G1S 47.19 (0.10) 21.79 (0.04) 33.53 (0.03)
@@ C.3 HetLoRA
C.4 Is the standard load balancing loss sufficient?
The standard load balancing loss encourages equal assignment of tokens to each expert. When the number of experts gets larger, there might not be enough tokens routed to the generalists, which might lead to a under-developed general knowledge. We will verify if this is indeed true.
To encourage enough tokens to be routed to the generalist expert such that more general knowledge can be developed, we modify our load-balancing loss by introducing importance weighting. As we separate the 0-th expert to be the generalist expert and conduct Top-2 routing, the modified load balancing loss is as follows:
@@ C.4 Is the standard load balancing loss sufficient?
$j$ is the expert index and $p(x)=[p_{j}(x)]_{j=1}^{n_{i}}$ is the logit output from the routing network for a specific token $x$. The idea is that one of the top 2 tokens should always be routed to the generalist expert, i.e. the 0-th expert. Thus, $\frac{p_{0}}{1/2}$ should be equal to $\frac{p_{i}}{1/2(n_{i}-1)}$ for $i\neq 0$. As the original load balancing loss encourages uniform distribution, this modification encourages the generalist expert to have a routing probability of 0.5 on expectation. Note that when $n_{i}=2$, this ${\mathcal{L}}_{i}^{\text{LB}}$ is the same as the original load balancing loss as proposed in Fedus et al. (2022a).
@@ Appendix D Additional Experiments
We replicate the experiments in Section 4.3 with the SlimPajama dataset, where we assign four times as many tokens to ArXiv User and Book User as to Stack Exchange User and Codes User.
More Specialists Help with Higher Data Quantity. From Figure 11, it is evident that ArXiv User and Book User, with abundant local data, benefit from having more local experts.
Figure 11: Test Perplexity during training for the SlimPajama setup. ArXiv User and Book User have more local data and thus benefit from having more experts. The numbers in the legend indicate the number of experts $n_{i}$ within each user. Top-2 routing is performed.
@@
Figure 12: In this SlimPajama setup, Stack Ex User and Codes User despite having low resources locally, overfit slightly on their small-sized local data. Numbers in the legend denote the number of experts $n_{i}$ within each user. Top2 routing is performed.
Specialists Can Benefit Generalists. Low-resourced users that can only support a single expert setup still benefit from collaboration, as the generalist knowledge is refined through a more detailed distinction between specialist and generalist roles via other high-resourced users. This is indicated by the enhanced performances for Stack Exchange and Codes Users.
Figure 13: In this SlimPajama setup, Stack Ex User and Codes User, despite having only one expert locally, still benefit from other users having more experts, thereby enhancing the generalist’s performance. The numbers in the legend indicate the number of experts, $n_{i}$, within each user. Top-2 routing is applied when $n_{i}$geq 2